OpenStudy (anonymous):

a machine takes 5.4 hours to make 9 parts. At that rate how many parts can the machine nake in 25.8 hours

OpenStudy (freckles):

set up \[\frac{9 parts}{5.4 hours}=\frac{x parts}{25.8 hours} \\ \frac{9}{5.4}=\frac{x}{25.8}\]

OpenStudy (freckles):

solve for x

OpenStudy (dan815):

25.8/5.4 * 9